X has the following PDF:
f(x)={Cxe^(-x/2) , x>0 and 0 , x=<0 }
a. Find the value of C.
b. Find the mean of X.
c. Find the 2nd Quartile of X.
d. Also, find the standard deviation of X.
1
Expert's answer
2020-04-01T16:22:59-0400
"f(x)= \begin{cases}
 Cxe^{-x/2} & x>0 \\
 0 & x\leq 0
\end{cases}"

a. Find the value of C. 


"1=\displaystyle\int_{-\infin}^\infin f(x)dx=\displaystyle\int_{0}^\infin Cxe^{-x/2}dx=""=\lim\limits_{A\rarr\infin}\bigg(\displaystyle\int_{0}^\infin Cxe^{-x/2}dx\bigg)"


"\int{xe^{-x/2}}dx=-2xe^{-x/2}+2\int{e^{-x/2}}dx=""=-2xe^{-x\/2}-4e^{-x\/2}+c_1"

"1=C\lim\limits_{A\rarr\infin}\big[-2xe^{-x/2}-4e^{-x/2}\big]\begin{matrix}
 A\\
 0
\end{matrix}="

"=C(-0-0+0+4)=4C=>C={1\over 4}"

"f(x)= \begin{cases}
 {1\over 4}xe^{-x/2} & x>0 \\
 0 & x\leq 0
\end{cases}"

b. Find the mean of X. 


"mean=E(X)=\displaystyle\int_{-\infin}^\infin xf(x)dx=""=\displaystyle\int_{0}^\infin {1\over 4}x^2 e^{-x/2}dx"

"\int{x^2e^{-x/2}}dx=-2x^2 e^{-x/2}+4\int{xe^{-x/2}}dx=""=-2x^2 e^{-x\/2}-8xe^{-x\/2}-16e^{-x\/2}+c_2"

"mean=E(X)=""={1\over 4}\lim\limits_{A\rarr\infin}\big[-2x^2 e^{-x/2}-8xe^{-x/2}-16e^{-x/2}\big]\begin{matrix}
 A\\
 0
\end{matrix}=""={1\over 4}(-0-0-0+0+0+16)=4"

c. Find the 2nd Quartile of X. 


"F(x)=\displaystyle\int_{-\infin}^x f(t)dt"

"\displaystyle\int_{0}^x {1\over 4}te^{-t/2}dt={1\over 4}\big[-2te^{-t/2}-4e^{-t/2}\big]\begin{matrix}
 x\\
 0
\end{matrix}="

"={1\over 4}(-2xe^{-x/2}-4e^{-x/2}+0+4)"

"F(0)=0"

Then


"F(x)= \begin{cases}
 -{1\over 2}xe^{-x/2}-e^{-x/2}+1 & x>0 \\
 0 & x\leq 0
\end{cases}"

"F(x_m)={1\over 2}=> -{1\over 2}x_me^{-x_m/2}-e^{-x_m/2}+1={1\over 2}=>"

"=> (x_m+2)e^{-x_m\/2}=1"



"median=2^{nd}Quartile=x_m\approx3.357"


d. Also, find the standard deviation of X.


"Var(X)=\sigma^2=E(X^2)-[E(X)]^2"

"E(X^2)=\displaystyle\int_{-\infin}^\infin x^2f(x)dx="

"=\displaystyle\int_{0}^\infin {1\over 4}x^3 e^{-x/2}dx"

"\int{x^3e^{-x/2}}dx=-2x^3 e^{-x/2}+6\int{x^2e^{-x/2}}dx="

"=-2x^3 e^{-x\/2}+6(-2x^2 e^{-x\/2}-8xe^{-x\/2}-16e^{-x\/2})="

"=-2x^3 e^{-x\/2}-12x^2 e^{-x\/2}-48xe^{-x\/2}-96e^{-x\/2}"

"E(X^2)={1\over 4}\lim\limits_{A\rarr\infin}\big[-2x^3 e^{-x/2}-12x^2 e^{-x/2}-48xe^{-x/2}-96e^{-x/2}\big]\begin{matrix}
 A\\
 0
\end{matrix}="

"={1\over 4}(-0-0-0+0+0+96)=24"


"Var(X)=\sigma^2=E(X^2)-[E(X)]^2=24-(4)^2=8"

"\sigma=\sqrt{8}=2\sqrt{2}\approx2.8284"